OpenStudy (anonymous):

I'm stuck at a simplifying-indices problem... [9^2 x 3^(n-1)]/[3^n + 3^(n-1)]

OpenStudy (anonymous):

The answer key on the book says 27/4, though. D: There's this rule that I don't understand. Oh, and the "x" is multiplication, btw... just noticed that.

OpenStudy (anonymous):

-_- for multiplication people usually use * for a multiplication not x

hero (hero):

There's no way you end up with 27/4 if that's your original problem.

hero (hero):

I get 81/4 btw

OpenStudy (anonymous):

@BVB_army, haha, sorry. Zombie-typing, I guess. :P @Hero, ... Really? How'd you get that? What I don't understand is how to take the common factor from the denominator. D:

hero (hero):

All you have to do is utilize rules of exponents to simplify

hero (hero):

\[3^{n-1}=3^n \times 3^{-1}\]

OpenStudy (anonymous):

... Uh, so then it becomes. 3^n + 3^n/3. Right? Then it should become 4(3^(n-1)), right? <-- That's the part I don't get... sorry to bother you with all these questions. D: Thank you, thank you.

hero (hero):

The steps for this is quite interesting. I'll have to show you on vyew.

OpenStudy (anonymous):

I'm not exactly sure what a Vyew is... but I'm making the account now.

OpenStudy (anonymous):

@Hero, thank you again. :D

hero (hero):

yw